# Infinite-time Mean Field FBSDEs and Viscosity Solutions to Elliptic Master Equations

### v2

@@ Abstract
This paper presents a further investigation of the properties of infinite-time mean field FBSDEs and elliptic master equations, which were introduced in [14] as mathematical tools for solving discounted infinite-time mean field games. By establishing the continuous dependence of the FBSDE solutions on their initial values, we prove the flow property of the mean field FBSDEs. Furthermore, we prove that, at the Nash equilibrium, the value function of the representative player constitutes a viscosity solution to the corresponding elliptic master equation. Our work extends the classical theory of finite-time mean field games and parabolic master equations to the infinite-time setting.
@@ 1 Introduction
The study of mean field games was initiated independently by Lasry-Lions (see [7, 8, 9]) and Huang-Malhamé-Caines [6], which is an analysis of limit models for symmetric weakly interacting $(N+1)$-player differential games. The parabolic master equation plays a crucial role in the analysis of mean field games, which was introduced by Lions in lectures [10]. It describes the strategic interaction between a representative player and the collective environment. We refer the reader to [4, 3, 5] for a comprehensive exposition on the subject.
Forward-backward stochastic differential equations (FBSDEs) also serve as a powerful tool for the study of mean field games. The investigation of general nonlinear BSDEs was pioneered by Pardoux and Peng [11, 12] in the early 1990s. [13] studied the infinite-time FBSDEs and established connections with quasilinear elliptic PDEs. Recently, [1] extended this framework to the Mckean-Vlasov FBSDEs. This paper establishes the existence and uniqueness theorem for a broader class of infinite-time FBSDEs, and employs it to address the existence of viscosity solutions for elliptic master equations.
In the recent work [14], we proposed the discounted infinite-time mean field game model and elliptic master equation, which extends the traditional framework to infinite-time case. Within this framework, we introduced the following two systems of infinite-time forward-backward stochastic differential equations (FBSDEs). The first one represents the state process of the social equilibrium, while the second denotes the state process of the representative player with initial state $x$.
@@
$H(x,\mu,y)\triangleq\min_{a\in\mathbb{R}}\left[b(x,\mu,a)\cdot y+f(x,\mu,a)\right],$ (1.3)
and $\hat{\alpha}(x,y)$ is the unique minimizer. Because we assume that $f(x,\mu,a)=f_{0}(x,\mu)+f_{1}(x,a)$ and $b(x,\mu,a)=b_{0}(x,\mu)+b_{1}(x,a)$, the value of minimizer $\hat{\alpha}$ is independent of $\mu$. After further assuming that $f$ and $b$ possess good smoothness properties, we can obtain the following relationship:
@@
In [14], we introduced the elliptic master equation:
@@
Here $\partial_{x},\partial_{xx}$ are standard spatial derivatives, $\partial_{\mu},\partial_{\tilde{x}\mu}$ are $W_{2}$-Wasserstein derivatives, $\tilde{\xi}$ is a random variable with law $\mu$ and $\tilde{\mathbb{E}}$ is the expectation with respect to its law. We have proved that if the master equation (1.5) admits a classical solution with sufficient regularity, we derive the following representation for Equation (1.1) and (1.2):
@@
$Y_{t}^{x}=\partial_{x}U(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}}),\quad Z_{t}^{x}=\partial_{xx}U(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}}).$ (1.7)
In this paper, we prove that (1.1) and (1.2) possess the flow property:
@@
is a viscosity solution to the master equation (1.5).
@@ 2 Preliminaries
We will use the filtered probability space $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$ endowed with a Brownian motion $B$. Its filtration $\mathbb{F}\triangleq(\mathcal{F}_{t})_{t\geq 0}$ is augmented by all $\mathbb{P}$-null sets and a sufficiently rich sub-$\sigma$-algebra $\mathcal{F}_{0}$ independent of $B$, such that it can support any probability measure on $\mathbb{R}$ with finite second moment.
@@
${\cal W}_{p}(\mu,\nu):=\inf\Big\{\big(\mathbb{E}[|\xi-\eta|^{q}]\big)^{1/q}:~\mbox{for all $\xi\in\mathbb{L}^{p}({\cal F};\mu)$, $\eta\in\mathbb{L}^{p}({\cal F};\nu)$}\Big\}.$
@@
To simplify, we set $L^{2}_{K}\triangleq L^{2}_{K}(0,\infty,\mathbb{R})$.
@@
Let ${\cal C}^{0}({\cal P}_{2})$ denote the set of ${\cal W}_{2}$-continuous functions $U:{\cal P}_{2}\to\mathbb{R}$. For ${\cal C}^{1}({\cal P}_{2})$, we mean the space of functions $U\in{\cal C}^{0}({\cal P}_{2})$ such that $\partial_{\mu}U$ exists and is continuous on ${\cal P}_{2}\times\mathbb{R}$, which is uniquely determined by (2.3). Let ${\cal C}^{2,1}(\mathbb{R}\times{\cal P}_{2})$ denote the set of continuous functions $U:\mathbb{R}\times{\cal P}_{2}\to\mathbb{R}$ such that $\partial_{x}U,\partial_{xx}U$ exist and are joint continuous on $\mathbb{R}\times{\cal P}_{2}$, $\partial_{\mu}U,\partial_{x\mu}U,\partial_{\tilde{x}\mu}U$ exist and are continuous on $\mathbb{R}\times{\cal P}_{2}\times\mathbb{R}$. Let ${\cal C}^{3,1}(\mathbb{R}\times{\cal P}_{2})$ denote the set of continuous functions $U:\mathbb{R}\times{\cal P}_{2}\to\mathbb{R}$ such that $\partial_{x}U,\partial_{xx}U,\partial_{xxx}U$ exist and are joint continuous on $\mathbb{R}\times{\cal P}_{2}$, $\partial_{\mu}U,\partial_{x\mu}U,\partial_{\tilde{x}\mu}U,\partial_{xx\mu}U,\partial_{x\tilde{x}\mu}U$ exist and are continuous on $\mathbb{R}\times{\cal P}_{2}\times\mathbb{R}$.
@@
Finally, we consider the space $\Theta\triangleq[0,+\infty)\times\mathbb{R}\times{\cal P}_{2}$, and let ${\cal C}^{1,2,1}(\Theta)$ denote the set of continuous functions $U:\Theta\to\mathbb{R}$ which has the following continuous derivatives: $\partial_{t}U$, $\partial_{x}U$, $\partial_{xx}U$, $\partial_{\mu}U$, $\partial_{x\mu}U$, $\partial_{\tilde{x}\mu}U.$ One crucial property of functions $U\in{\cal C}^{1,2,1}(\Theta)$ is the Itô’s formula [2, 4]. For $i=1,2$, let $\mathrm{d}X^{i}_{t}\triangleq b^{i}_{t}\mathrm{d}t+\sigma^{i}_{t}\mathrm{d}B_{t},$ where $b^{i}:[0,+\infty)\times\Omega\to\mathbb{R}$ and $\sigma^{i}:[0,+\infty)\times\Omega\to\mathbb{R}$ are $\mathbb{F}$-progressively measurable and bounded (for simplicity), and $\rho_{t}\triangleq{\cal L}_{X^{2}_{t}}$. Fix $T>0$ and let all conditions be restricted to the interval [0,T]. Then we have
$\begin{split}\mathrm{d}U(t,X^{1}_{t},\rho_{t})=&\Big[\partial_{t}U+\partial_{x}U\cdot b^{1}_{t}+\frac{1}{2}\partial_{xx}U[\sigma_{t}^{1}]^{2}\Big](t,X^{1}_{t},\rho_{t})\mathrm{d}t\\ &+\Big(\tilde{\mathbb{E}}_{{\cal F}_{t}}\big[\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})(\tilde{b}^{2}_{t})+\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})[\tilde{\sigma}_{t}^{2}]^{2}\big]\Big)\mathrm{d}t\\ &+\partial_{x}U(t,X^{1}_{t},\rho_{t})\sigma_{t}^{1}\mathrm{d}B_{t}.\end{split}$ (2.4)
Here $\tilde{\mathbb{E}}_{{\cal F}_{t}}$ is the conditional expectations given ${\cal F}_{t}$ corresponding to the probability measure $\tilde{\mathbb{P}}$ .
3 Solutions to infinite-time FBSDEs
For the needs of subsequent problems, we aim to establish a more general theorem on the existence and uniqueness of solutions for infinite-time FBSDEs. Consider the following form of infinite-time FBSDEs:
$\begin{cases}\mathrm{d}X_{t}=G(t,\omega,X_{t},Y_{t},{\cal L}_{X_{t}})\mathrm{d}t+\sigma\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}=-F(t,\omega,X_{t},Y_{t},\mathcal{L}_{X_{t}})\mathrm{d}t+Z_{t}\mathrm{d}B_{t},\\ X_{0}=\xi,\end{cases}$ (3.1)
where $G,F:\mathbb{R}_{+}\times\Omega\times\mathbb{R}\times\mathbb{R}\times{\cal P}_{2}\rightarrow\mathbb{R}$ are two progressively measurable functions, $\sigma\in\mathbb{R}$ is a constant and $\xi$ is an ${\cal F}_{0}$-measurable square integrable random variable. For any $(v_{t})\in L_{K}^{2}$, we define the exponentially weighted $L^{2}$ norm
For simplicity, we only solve (3.1) for one dimensional $(X_{t},Y_{t},Z_{t})$ and starting time $t_{0}=0$ , but our result can be easily generalized to multidimensional case and arbitrary starting time $t_{0}>0$. The key idea of our proof follows [1, 13].
Assumption 3.1
Assume that for some constant $K$, the functions $F$ and $G$ satisfy:
(i) For any $L_{K}^{2}$ processes $(X_{t},Y_{t})$, $G(t,\omega,X_{t},Y_{t},{\cal L}_{X_{t}})$ and $F(t,\omega,X_{t},Y_{t},{\cal L}_{X_{t}})$ belong to $L_{K}^{2}$.
(ii) There exists a positive constant $\ell$ such that for any $x,x^{\prime},y,y^{\prime}\in\mathbb{R},\mu,\mu^{\prime}\in{\cal P}_{2}$
$|G(t,\omega,x,y,\mu)-G(t,\omega,x^{\prime},y^{\prime},\mu^{\prime})|+|F(t,\omega,x,y,\mu)-F(t,\omega,x^{\prime},y^{\prime},\mu^{\prime})|$ (3.3)
(iii) There exists a constant $\kappa>K/2$, such that for any $t\geq 0$ and any square integrable random variables $X,X^{\prime},Y,Y^{\prime}$,
$\mathbb{E}\left[-K\hat{X}\hat{Y}-\hat{X}(F(t,\omega,U)-F(t,\omega,U^{\prime}))+\hat{Y}(G(t,\omega,U)-G(t,\omega,U^{\prime}))\right]$ (3.4)
$\leq-\kappa\mathbb{E}\left[\hat{X}^{2}+\hat{Y}^{2}\right],$
where $\hat{X}\triangleq X-X^{\prime},\hat{Y}\triangleq Y-Y^{\prime}$ and $U\triangleq(X,Y,\mathcal{L}_{X}),U^{\prime}\triangleq(X^{\prime},Y^{\prime},\mathcal{L}_{X^{\prime}}).$
Theorem 3.2
Under Assumption 3.1, for each ${\cal F}_{0}$-measurable square integrable random variable $\xi$ , (3.1) has a unique solution $(X_{t},Y_{t},Z_{t})$ in $L_{K}^{2}$.
Proof.  First, we prove the uniqueness. Suppose there exist two solutions $(X_{t},Y_{t},Z_{t})$, $(X_{t}^{\prime},Y_{t}^{\prime},Z_{t}^{\prime})$ in $L_{K}^{2}$ to (3.1), and denote
$\hat{X}\triangleq X-X^{\prime}\quad\hat{Y}\triangleq Y-Y^{\prime}\quad\hat{Z}\triangleq Z-Z^{\prime}.$ (3.5)
We choose a sequence of $T_{i}\to\infty$ such that
$\mathbb{E}\left[\mathrm{e}^{-KT_{i}}\hat{X}_{T_{i}}\hat{Y}_{T_{i}}\right]\rightarrow 0.$ (3.6)
Applying Itô’s formula to $\mathrm{e}^{-Kt}\hat{X}_{t}\hat{Y}_{t}$, we get that
$\mathbb{E}\left[\mathrm{e}^{-KT_{i}}\hat{X}_{T_{i}}\hat{Y}_{T_{i}}\right]$ (3.7)

### v3

@@ Abstract
This paper presents a further investigation of the properties of infinite-time mean field FBSDEs and elliptic master equations, which were introduced in [14] as mathematical tools for solving discounted infinite-time mean field games. By establishing the continuous dependence of the FBSDE solutions on their initial values, we prove the flow property of the mean field FBSDEs. And then, we prove that, at the Nash equilibrium, the value function of the representative player constitutes a viscosity solution to the corresponding elliptic master equation. Furthermore, we establish the uniqueness of classical solutions to the elliptic master equation under displacement monotonicity and certain growth assumptions.
@@ 1 Introduction
The study of mean field games was initiated independently by Lasry-Lions [7, 8, 9] and Huang-Malhamé-Caines [6], which is an analysis of limit models for symmetric weakly interacting $(N+1)$-player differential games. The parabolic master equation plays a crucial role in the analysis of mean field games, which was introduced by Lions in lectures [10]. It describes the strategic interaction between a representative player and the collective environment. We refer the reader to [4, 3, 5] for a comprehensive exposition on the subject.
Forward-backward stochastic differential equations (FBSDEs) also serve as a powerful tool for the study of mean field games. The theory of general nonlinear BSDEs was pioneered by Pardoux and Peng [11, 12] in the early 1990s. [13] studied the infinite-time FBSDEs and established connections with quasilinear elliptic PDEs. Recently, [1] extended this framework to the McKean-Vlasov FBSDEs. This paper investigates the properties of infinite-time mean field FBSDEs, and employs them to address the existence of viscosity solutions for elliptic master equations.
In the recent work [14], we proposed the discounted infinite-time mean field games and elliptic master equations, which extends the traditional framework to infinite-time case. In our framework, a representative player interacts with a continuum of other players (also referred to as the population or social equilibrium). Let $\mu_{0}\in{\cal P}(\mathbb{R})$ be the initial state distribution in the society. For the player($x$), a representative player with initial state $x$, the dynamic of its state is given by
$X^{x,\beta}_{t}=x+\int_{0}^{t}b(X^{x,\beta}_{t},\mu_{s},\beta_{s})\mathrm{d}s+B_{t},$ (1.1)
where $\beta$, an $\mathbb{R}$-valued progressively measurable stochastic process, is the strategy of the player($x$) and $(\mu_{t})_{t\geq 0}$ is the population distribution flow start at $\mu_{0}$. Denote by $\beta^{x}$ the optimal strategy of the representative player($x$) derived from the stochastic optimization problem:
$\min_{\beta}J^{\mu}(\beta)\triangleq\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}f(X_{t}^{x,\beta},\mu_{t},\beta_{t})\mathrm{d}t\right].$ (1.2)
Since the population consists of a multitude of individuals, the macroscopic distribution should satisfy
$\mu_{t}(\cdot)=\int\mathbb{P}(X_{t}^{x,\beta^{x}}\in\cdot)\mu_{0}(\mathrm{d}x).$ (1.3)
To solve this fixed-point problem, we introduce the following two systems of FBSDEs. The first one represents the state process of the social equilibrium, while the second denotes the state process of the representative player with initial state $x$.
$\begin{cases}\mathrm{d}X_{t}^{\xi}=\partial_{y}H(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{\xi})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\left[\partial_{x}{H}(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{\xi})-rY_{t}^{\xi}\right]\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (1.4)
@@
$H(x,\mu,y)\triangleq\min_{a\in\mathbb{R}}\left[b(x,\mu,a)\cdot y+f(x,\mu,a)\right].$ (1.6)
For convenience, we assume that $f(x,\mu,a)=f_{0}(x,\mu)+f_{1}(x,a),b(x,\mu,a)=b_{0}(x,\mu)+b_{1}(x,a)$, and denote by $\hat{\alpha}(x,y)$ the unique minimizer of (1.6). After further assuming that $f$ and $b$ possess good smoothness properties, we can obtain the following relationship:
@@
In [14], we define $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ with ${\cal L}_{\xi}=\mu$ and demonstrate that $\mathcal{V}$ serves as a viscosity solution to the distribution-dependent PDE below:
$\begin{split}r\mathcal{U}(x,\mu)=&\partial_{x}H(x,\mu,\mathcal{U}(x,\mu))+\partial_{y}H(x,\mu,\mathcal{U}(x,\mu))\cdot\partial_{x}\mathcal{U}(x,\mu)+\frac{1}{2}\partial_{xx}\mathcal{U}(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\mathcal{U}(x,\mu,\tilde{\xi})+\partial_{\mu}\mathcal{U}(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\mathcal{U}(\tilde{\xi},\mu))\right].\end{split}$ (1.8)
Here $\partial_{x},\partial_{xx}$ are standard spatial derivatives, $\partial_{\mu},\partial_{\tilde{x}\mu}$ are ${\cal W}_{2}$-Wasserstein derivatives, $\tilde{\xi}$ is a random variable with law $\mu$ and $\tilde{\mathbb{E}}$ is the expectation with respect to its law. We also introduce the elliptic master equation:
@@
We have proved that if this master equation admits a classical solution with sufficient regularity, we can derive the following representation for FBSDEs (1.4) and (1.5):
@@
$Y_{t}^{x,\xi}=\partial_{x}U(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}}),\quad Z_{t}^{x,\xi}=\partial_{xx}U(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}}).$ (1.11)
In this paper, we make above relation rigorous and proceed to study the properties of the elliptic master equation. We prove that solutions to FBSDEs (1.4) and (1.5) possess the flow property:
@@
is a viscosity solution to elliptic master equation (1.9), and satisfies the relationship
$\partial_{x}V(x,\mu)=Y_{0}^{x,\xi}.$ (1.15)
Finally, we address the uniqueness of classical solutions to the elliptic master equation. We introduce a displacement monotonicity condition (see [5]), which states that for any square integrable random variables $\xi_{1},\xi_{2}$, the following inequality holds:
$\mathbb{E}\left[(\xi_{1}-\xi_{2})\left(\partial_{x}U(\xi_{1},\mathcal{L}_{\xi_{1}})-\partial_{x}U(\xi_{2},\mathcal{L}_{\xi_{2}})\right)\right]\geq 0.$ (1.16)
Under suitable growth conditions and this displacement monotonicity condition, we establish the uniqueness of classical solutions to the elliptic master equation.
@@ 2 Preliminaries
We will use a filtered probability space $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$ endowed with a Brownian motion $B$. Its filtration $\mathbb{F}\triangleq(\mathcal{F}_{t})_{t\geq 0}$ is augmented by all $\mathbb{P}$-null sets and a sufficiently rich sub-$\sigma$-algebra $\mathcal{F}_{0}$ independent of $B$, such that it can support any measure on $\mathbb{R}$ with finite second moment.
@@
${\cal W}_{p}(\mu,\nu)\triangleq\inf\Big\{\big(\mathbb{E}[|\xi-\eta|^{q}]\big)^{1/q}:\mbox{for all $\xi\in\mathbb{L}^{p}({\cal F};\mu)$, $\eta\in\mathbb{L}^{p}({\cal F};\nu)$}\Big\}.$
@@
To simplify, we set $L^{2}_{K}\triangleq L^{2}_{K}(0,\infty,\mathbb{R})$ and define the exponentially weighted $L^{2}$ norm
For each $\mathcal{F}_{0}$-measurable square integrable random variable $\xi$ , we consider the following infinite-time FBSDE:
$\begin{cases}\mathrm{d}X_{t}=G(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}=-F(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+Z_{t}\mathrm{d}B_{t},\\ X_{0}=\xi.\end{cases}$ (2.4)
Here $G,F:\mathbb{R}^{+}\times\mathbb{R}^{2}\times{\cal P}_{2}\to\mathbb{R}$ are two measurable functions, and satisfy the following assumptions:
Assumption 2.1
(i) There exists a positive constant $\ell$ such that for any $x,x^{\prime},y,y^{\prime}\in\mathbb{R},\mu,\mu^{\prime}\in{\cal P}_{2}$
$|G(t,x,y,\mu)$ $-G(t,x^{\prime},y^{\prime},\mu^{\prime})|+|F(t,x,y,\mu)-F(t,x^{\prime},y^{\prime},\mu^{\prime})|$ (2.5)
(ii) There exist constants $0<K<2\kappa$ such that for any $t\geq 0$ and any square integrable random variables $X,X^{\prime},Y,Y^{\prime}$
$\mathbb{E}\left[-K\hat{X}\hat{Y}-\hat{X}(F(t,U)-F(t,U^{\prime}))+\hat{Y}(B(t,U)-B(t,U^{\prime}))\right]$ (2.6)
$\leq-\kappa\mathbb{E}\left[\hat{X}^{2}+\hat{Y}^{2}\right],$
where $\hat{X}=X-X^{\prime},\hat{Y}=Y-Y^{\prime}$ and $U=(X,Y,\mathcal{L}_{X}),U^{\prime}=(X^{\prime},Y^{\prime},\mathcal{L}_{X^{\prime}}).$
The following lemma states the existence and uniqueness of a solution to FBSDE(2.4). For the detailed proof, we refer the reader to ([1], Theorem 2.1).
Lemma 2.2
Under Assumption 2.1, the FBSDE (2.4) admits a unique solution in $L_{K}^{2}$.
@@
Let $C^{0}({\cal P}_{2})$ denote the set of ${\cal W}_{2}$-continuous functions $U:{\cal P}_{2}\to\mathbb{R}$. For $C^{1}({\cal P}_{2})$, we mean the space of functions $U\in C^{0}({\cal P}_{2})$ such that $\partial_{\mu}U$ exists and is continuous on ${\cal P}_{2}\times\mathbb{R}$, which is uniquely determined by (2.7). Let $C^{2,1}(\mathbb{R}\times{\cal P}_{2})$ denote the set of continuous functions $U:\mathbb{R}\times{\cal P}_{2}\to\mathbb{R}$ such that $\partial_{x}U,\partial_{xx}U$ exist and are joint continuous on $\mathbb{R}\times{\cal P}_{2}$, $\partial_{\mu}U,\partial_{x\mu}U,\partial_{\tilde{x}\mu}U$ exist and are continuous on $\mathbb{R}\times{\cal P}_{2}\times\mathbb{R}$.
@@
Finally, we consider the space $\Theta\triangleq[0,T]\times\mathbb{R}\times{\cal P}_{2}$ for some $T>0$, and let $C^{1,2,1}(\Theta)$ denote the set of continuous functions $U:\Theta\to\mathbb{R}$ which has the following continuous derivatives: $\partial_{t}U$, $\partial_{x}U$, $\partial_{xx}U$, $\partial_{\mu}U$, $\partial_{x\mu}U$, $\partial_{\tilde{x}\mu}U.$ One crucial property of functions $U\in C^{1,2,1}(\Theta)$ is the Itô’s formula[2, 4]. For $i=1,2$, let $\mathrm{d}X^{i}_{t}\triangleq b^{i}_{t}\mathrm{d}t+\sigma^{i}_{t}\mathrm{d}B_{t},$ where $b^{i}:[0,T]\times\Omega\to\mathbb{R}$ and $\sigma^{i}:[0,T]\times\Omega\to\mathbb{R}$ are $\mathbb{F}$-progressively measurable and bounded (for simplicity), and $\rho_{t}\triangleq{\cal L}_{X^{2}_{t}}$. Suppose that for every compact subset ${\cal K}\subset\mathbb{R}\times{\cal P}_{2}$, it holds that: